OpenStudy (wildemo):

The function f(x) = −x2 + 48x − 432 models the profit, in dollars, a game developer makes for selling a new game, where x is the number of games sold and f(x) is the amount of profit. Part A: Determine the vertex. What does this calculation mean in the context of the problem? what i've got so far is y=-1(x^2-48+576-576+432) y=-1[(x-24)^2-144). I believe I am right and am confused about the next part o the problem. I will award medals.

OpenStudy (wildemo):

Part B: Determine the x-intercepts. What do these values mean in the context of the problem? I have gotten that the answer is (12,0) and (36,0) but i do no9t know what the represent in the context of the problem.

ganeshie8 (ganeshie8):

f(x) = −x2 + 48x − 432 = -1(x - 24)^2 + 144 so the vertex would be (24, 144) right ?

OpenStudy (wildemo):

Yes, that is what I was thinking but wasn't sure if it was correct

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

As you can see, a parabola has only one vertex point.

ganeshie8 (ganeshie8):

Notice that vertex is the top of the hill. The maximum point.

ganeshie8 (ganeshie8):

In our problem, what do x and f(x) represent ?

OpenStudy (wildemo):

x is the number of games sold and f(x) is the amount of profit

OpenStudy (wildemo):

f(x)=y

ganeshie8 (ganeshie8):

Since the maximum value occurs at the vertex, what does the point (24, 144) represent here ?

OpenStudy (wildemo):

24 is the max amount of games sold. 144 is the amount of profit.

ganeshie8 (ganeshie8):

Can we say that the maximum value of the profit is 144, and it occurs when 24 games are sold ?

OpenStudy (wildemo):

yes i agree lol

ganeshie8 (ganeshie8):

24 is the number of games to be sold to achieve the maximum profit 144 is the maximum profit

ganeshie8 (ganeshie8):

Good :)

OpenStudy (wildemo):

Im still little confused about part b though :/

OpenStudy (wildemo):

@ganeshie8 ^^^